OpenStudy (anonymous):

If y= -1/2x +1, what would be the the x intercept?

ganeshie8 (ganeshie8):

to find x-intercept, replace y wid 0 and solve x

ganeshie8 (ganeshie8):

y= -1/2x +1 0 = -1/2x + 1 solve x

OpenStudy (anonymous):

1

ganeshie8 (ganeshie8):

try agaiin

OpenStudy (anonymous):

Oh it's 2

ganeshie8 (ganeshie8):

0 = -1/2x + 1 1/2x = 1 x = 2

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

Now I get it, thanks for the help.

ganeshie8 (ganeshie8):

glad to hear... u wlc :)